The result (in hexadecimal) of adding the two hexadecimal numbers BABA and A9 is ______ (Use only uppercase letters for hex digits)?
Answer #1

Carry:    011
=============
         BABA 
         + A9   


A + 9 = 19 in decimal, carry = 1, result = 3(Hex)
1 + B + A = 22 in decimal, carry = 1, result = 6(Hex)
1 + A = 11 in decimal, carry = 0, result = B(Hex)
0 + B = 12 in decimal, carry = 0, result = B(Hex)

Answer = BB63

**************************************************

Thanks for your question. We try our best to help you with detailed answers, But in any case, if you need any modification or have a query/issue with respect to above answer, Please ask that in the comment section. We will surely try to address your query ASAP and resolve the issue.

Please consider providing a thumbs up to this question if it helps you. by Doing that, You will help other students, who are facing similar issue.